OpenStudy (anonymous):

what is the y/x intercept of y=3x+3? PLEASE PLEASE PLEASE HELP!!!

OpenStudy (anonymous):

for y-intercept, you know x would be zero. sub in zero for x to get y=3(0)+3 which gives y=3 as y-int for x-int, sub in 0 for y and solve for x. you'll get 0=3x+3 -3=3x x=-1 is your x-int